# GroundCap: A Visually Grounded Image Captioning Dataset

### v2

@@ 5.1 Training Methodology
We fine-tuned Pixtral-12B using Low-Rank Adaptation (LoRA) [32] with a rank of 16 and alpha scaling factor of 32, targeting both self-attention layers (query, key, value, output projections) and MLP layers (gate, up, down projections) of the language components. This approach reduces memory requirements while maintaining model quality.
@@ 5.2 Evaluation
Evaluation of the Pixtral-trained model was conducted on our test set of 10,000 images, comparing performance between the model fine-tuned on automatic captions (Pixtral auto) and after additional fine-tuning on human-refined captions (Pixtral human).
@@
Table 1: Evaluation results on the GroundCap test set. Pixtral (auto) refers to the model after fine-tuning on automatically generated captions, while Pixtral (human) indicates performance after further fine-tuning on human-refined captions. We report precision (P), recall (R), F1, BLEU-4 (B-4), METEOR (M), CIDEr (C), SPICE (S), ROUGE-L, and gMETEOR (gM).
@@
Our results in Table 1 show that both versions of the model achieve strong grounding recall (0.95 and 0.96 respectively) but lower precision (0.61 and 0.58), resulting in F1 scores around 0.70. This indicates the models successfully reference most detected objects but occasionally include incorrect or imprecise groundings.
The language quality metrics show modest but consistent performance across both models. BLEU-4 scores of 0.18 and 0.19 reflect reasonable n-gram overlap with reference captions, while METEOR scores around 0.24 indicate acceptable semantic similarity. The improvement in CIDEr score from 0.46 to 0.51 after human fine-tuning suggests better alignment with consensus descriptions.
Our proposed gMETEOR metric, combining grounding accuracy with language quality, yields identical scores of 0.35 for both models. This suggests that while fine-tuning on human-refined captions produces some improvements in caption quality metrics, the trade-off between precision and recall in grounding accuracy results in similar overall performance.
These results establish a strong baseline for GroundCap while highlighting areas for potential improvement, particularly regarding grounding precision and semantic alignment with reference captions.
The comparable performance between automatically trained and human-refined models suggests that our automatic caption generation process produces reasonable training data, though there remains room for improvement in both grounding and language generation capabilities.
@@ 6 Human Evaluation
We conducted a comprehensive human evaluation to assess the quality of our grounded captioning system and compare it with baseline approaches. This evaluation involved multiple annotators rating different types of captions across several criteria, complemented by an analysis of automatic metrics and an exploration of using ChatGPT-4o for automated assessment.
@@ 6.1 Evaluation Setup
We compared three caption versions for each image: automatically generated captions from our pipeline, human-refined versions of these captions, and outputs from our fine-tuned Pixtral model. Each caption was evaluated by three different annotators, resulting in 2,274 total evaluations. To prevent bias, captions were presented to evaluators in random order and evaluators could not evaluate captions they had previously annotated. Evaluators used an interface showing the image with detected objects alongside the caption being assessed with highlighted grounding tags as shown in fig. 5.
@@ 6.2 Evaluation Criteria
6.3 Human Assessment Results
Table 2 presents the human evaluation scores across all criteria. Human-refined captions achieved the highest overall quality (4.34), outperforming both automatic captions (4.07) and Pixtral-generated ones (4.22). All caption types received high scores for language quality ($>4.8$), demonstrating Pixtral abilities at language generation.
Human-refined captions showed particular strength in grounding precision (4.23) and recall (4.33), indicating that human refinement improves the accuracy of object references while maintaining comprehensive coverage. Pixtral-generated captions achieved comparable grounding precision (4.22) to human-refined ones, demonstrating the model’s ability to learn accurate grounding from the smaller set of human-refined captions.
@@ 8 Conclusions and Future Work
This paper introduced GroundCap, a novel dataset for grounded captioning that provides detailed descriptions of visual scenes grounded on detected objects, actions, and locations using an unified grounding framework that maintains object identity across multiple references. The proposed gMETEOR which combines language quality with grounding accuracy, providing improved quality assessment. Additionally a baseline model for grounded captioning on GroundCap and conducted a human evaluation study to assess the quality of our automatically generated, baseline model, and human-refined captions. Our baseline model achieves strong grounding recall (0.96) while maintaining good language quality (METEOR score of 0.23). Human evaluation demonstrates the effectiveness of our approach, with human-refined captions receiving high ratings across all criteria (4.34/5.0 overall quality). Manual inspection revealed that the lowest-performing captions are often frames where few or no objects were detected or the image contents were unclear (e.g. in blurry frames). We found that ChatGPT-4o evaluations strongly correlate with human judgments (0.78 Pearson correlation for overall quality), suggesting its potential as a scalable evaluation tool, especially given the poor correlation between traditional metrics and human assessment. These results establish GroundCap as a valuable resource for advancing research in visually grounded language generation.

### v3

@@ 5.1 Training Methodology
We fine-tuned both models using Low-Rank Adaptation (LoRA) [32] with a rank of 16 and alpha scaling factor of 32, targeting both self-attention layers (query, key, value, output projections) and MLP layers (gate, up, down projections) of the language components. This approach reduces memory requirements while maintaining model quality.
@@ 5.2 Evaluation
Evaluation of both models was conducted on our test set of 10,000 images, comparing performance between the model fine-tuned on automatic captions (Pixtral auto / Qwen auto) and after additional fine-tuning on human-refined captions (Pixtral human / Qwen human).
@@
Table 2: Evaluation results on the GroundCap test set. Pixtral (auto) and Qwen (auto) refers to the models after fine-tuning on automatically generated captions, while Pixtral (human) and Qwen (human) indicates performance after further fine-tuning on human-refined captions. We report precision (P), recall (R), F1, BLEU-4 (B-4), METEOR (M), CIDEr (C), SPICE (S), ROUGE-L, and gMETEOR (gM).
@@
Qwen (auto) 0.53 0.29 0.28 0.12 0.19 0.14 0.22 0.29 0.19
Qwen (human) 0.66 0.25 0.21 0.07 0.13 0.05 0.12 0.19 0.13
Our results in Table 2 show clear performance differences between the two models. Pixtral outperforms Qwen across most metrics, achieving high grounding recall (0.95-0.96) with moderate precision (0.58-0.61), resulting in F1 scores around 0.70. Qwen shows the opposite pattern: lower recall (0.25-0.29) than precision (0.53-0.66), yielding lower F1 scores (0.21-0.28). Pixtral references most detected objects but includes some incorrect groundings, while Qwen is more selective but misses many objects. Pixtral’s better language quality was expected since the original training captions were generated using Pixtral’s policy. It achieves BLEU-4 scores of 0.18-0.19 and METEOR scores around 0.24, while Qwen scores substantially lower across all language metrics. Pixtral’s higher recall does not come from this policy advantage since both models use the same object detections generated by Mask2Former. This difference reflects how the models process visual-textual correspondences.
The comparable performance between automatically trained and human-refined models suggests that our automatic caption generation process produces reasonable training data, though there remains room for improvement in both grounding and language generation capabilities. These results establish Pixtral as the stronger baseline for GroundCap.
@@ 6 Human Evaluation
We conducted human evaluation to assess the quality of our grounded captioning system and compare it with baseline approaches. This evaluation involved multiple annotators rating different types of captions across several criteria, complemented by an analysis of automatic metrics and an exploration of using ChatGPT-4o for automated assessment. We focused our human evaluation on Pixtral only due to the costs of conducting human assessment across multiple models and because it yielded the best results in our automatic evaluation.
@@ 6.1 Evaluation Setup
We compared three caption versions for each image: automatically generated captions from our pipeline, human-refined versions of these captions, and outputs from the fine-tuned Pixtral model. Each caption was evaluated by three different annotators, resulting in 2,274 total evaluations. To prevent bias, captions were presented to evaluators in random order and evaluators could not evaluate captions they had previously annotated. Evaluators used an interface showing the image with detected objects alongside the caption being assessed with highlighted grounding tags as shown in fig. 5.
@@ 6.2 Human Assessment
Table 3 presents the human evaluation scores across all criteria. Human-refined captions achieved the highest overall quality (4.34), outperforming both automatic captions (4.07) and Pixtral-generated ones (4.22). All caption types received high scores for language quality ($>4.8$), demonstrating Pixtral abilities at language generation. Human-refined captions showed particular strength in grounding precision (4.23) and recall (4.33), indicating that human refinement improves the accuracy of object references while maintaining comprehensive coverage. Pixtral-generated captions achieved comparable grounding precision (4.22) to human-refined ones, demonstrating the model’s ability to learn accurate grounding from the smaller set of human-refined captions.
@@ 8 Conclusions and Future Work
This paper introduced GroundCap, a novel dataset for grounded captioning that provides detailed descriptions of visual scenes grounded on detected objects, actions, and locations using an unified grounding framework that maintains object identity across multiple references. The proposed gMETEOR which combines language quality with grounding accuracy, providing improved quality assessment. Additionally, a baseline model for grounded captioning on GroundCap and conducted a human evaluation study to assess the quality of our automatically generated, baseline model, and human-refined captions. Our baseline model achieves strong grounding recall (0.96) while maintaining good language quality (METEOR score of 0.23). Human evaluation demonstrates the effectiveness of our approach, with human-refined captions receiving high ratings across all criteria (4.34/5.0 overall quality). Manual inspection revealed that the lowest-performing captions are often frames where few or no objects were detected or the image contents were unclear (e.g., in blurry frames). We found that ChatGPT-4o evaluations strongly correlated with human judgments (0.78 Pearson correlation for overall quality), suggesting its potential as a scalable evaluation tool, especially given the poor correlation between traditional metrics and human assessment. These results establish GroundCap as a valuable resource for advancing research in visually grounded language generation.